OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line (4,-9);2x-5y=7

OpenStudy (anonymous):

Get Y free to on side.. And then you find the slope.. Use it for build a Equation of the given point..

OpenStudy (anonymous):

\[2x-5y=7\] \[y=(2x-7)\div5\] \[y=2x \div5-7\div5\] And the slope is then 2\5.. k=2.5 , and x =4 , y =-9 Use it : Y=kx+m

OpenStudy (anonymous):

will give....m = 53\5

OpenStudy (anonymous):

Equation is then... y=2.5x-53\5

OpenStudy (amistre64):

slope can be remembered as -a/b which in this equations is: -2/-5 = 2/5

OpenStudy (anonymous):

Sry forgott the "\"

OpenStudy (amistre64):

the short cut to a parallel line equation is to just plug in your values of (x,y) = get the "new" answer, for example: 2x-5y=7 ; (x=4,y=-9) 2(4) -5(-9) = ? 8 +45 = 53 "new" equation for PARALLEL line is: 2x -5y = 53 ;(4,-9)

OpenStudy (anonymous):

ahh. even shorter way thx... iam Used to work with y=kx+m .

OpenStudy (amistre64):

i used the y=mx+b as well, but after so many times, I noticed this pattern :)

OpenStudy (anonymous):

I am gonna go over! to Ax+by=C